OpenStudy (anonymous):

integral [ cos^4(x) sin^3(x) ] dx. show steps.

OpenStudy (ash2326):

\[\int \cos^{4} x\ \sin ^{3} x\ dx\] Let's rewrite this \[\int \cos^{4} x\ \sin x\ \sin ^{2} x\ dx\] we know that \(\sin ^2 x= 1-\cos^2 x\) , we get \[\int \cos^{4} x\ \sin x\ (1-\cos ^{2} x)\ dx\] Now let's substitute \[\cos x=t\] we get \[-\sin x\ dx=dt\] so our integral changes to \[\large \int t^4(1-t^2) -dt\] Can you take it from here? @alienbrain

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

i'm not sure, please continue.

OpenStudy (anonymous):

take a look at this http://tutorial.math.lamar.edu/pdf/Calculus_Cheat_Sheet_Integrals_Reduced.pdf

OpenStudy (anonymous):

wow, that's pretty cool!

OpenStudy (anonymous):

see...Products and (some) Quotients of Trig Functions first page !

OpenStudy (anonymous):

love the process breakdown

OpenStudy (anonymous):

ok, i should be good. thanks

OpenStudy (anonymous):

yw !